(a) z 0.81 for a right tail test for a difference in two proportions Round your answer to two decimal places. p-value the absolute tolerance is +/-0.01 SHOW HINT LINK TO TEXT (b) Z =-2.46 for a left tail test for a difference in two means Round your answer to three decimal places. p-value- the absolute tolerance is +/-0.001
Answer #1

a) P-value = P(Z > 0.81) = 1 - P(Z < 0.81) = 1 - 0.7910 = 0.2090

b) P-value = P(Z < -2.46) = 0.0069

c) P-value = 2 * P(Z > 2.3) = 2 * (1 - P(Z < 2.3)) = 2 * (1 - 0.9893) = 0.0214